[Jenne, Koji, Rana]
Thanks to turning off the AS55 analog whitening as well as the 1k:6k lead filter that Koji put into Darm's FM7, the DARM transition was more stable early in the evening.
The AS55 gain and offset did not change noticeably when we switched the AA on or off (switching happened while *not* using AS for any feedback). Earlier in the evening, we did also check what happened with PRMI and REFL33 AA on vs. off, and REFL33 did have a many tens of counts offset on both the I and Q input channels. I have turned the AA filters back on, but run LSCoffsets before trying to lock.
I'm not sure what was up, but somehow I couldn't lock the PRMI for about half an hour or so. Very frustrating. Eventually after futzing around, I was able to get it to lock with REFL33 in PRMI-only, and after that it worked again in PRFPMI with REFL165.
With FSS slow around 0.5, MC has been a bit fussy the last hour. Also frustrating.
Later on in the evening, I started taking out a bunch of the "sleep" commands from the up script, and many of the "press enter to continue" spots, but I think it might be moving too fast. That, or I'm just not catching where I have too much gain. Anyhow, near the middle/end of the CARM transition I am getting severe gain peaking at several hundred Hz. I think I need to use a lower final gain.
So, progress on DARM, but maybe a little more fine-tuning of CARM needed.
Here's a DARM loop measurement, taken after both CARM and DARM were RF-only:
DARM_RF-only_13Mar2015.pdf |